Charlene is the owner of Charlenes Computer Care (3C) and has been in the business of servicing computers and computer systems at affordable prices for just over five years. 3Cs primary clients are local businesses but include some residential customers. Their advertised approach is to analyze the problem, develop a range of IT options, explain their specific recommendation to the customer, and finally estimate the time and costs associated with the service prior to commencing work. Their approach has proven successful, but competition is getting much stiffer, and Charlene is concerned that her approach to determining costs and prices for any particular job is not correctly incorporating all of the companys costs. She has heard about the concept of cost allocation and asks her accountant to create an allocation system for the business. Her accountant decides to implement a single-driver allocation system and needs to develop a reliable cost function to estimate total indirect costs. She will use this estimate for the numerator of the single driver overhead rate, and she will use service hours as the cost driver and denominator in the overhead rate. To estimate the parameters of this cost function, the accountant first studied the behavior of last months costs, when the company spent 309 service hours repairing 27 computer systems. Here are the results of her analysis: Cost Item Utilities Office rent Administratiorn Supplies Training Fixed Cost $271 $3,293 $3,129 $0 $339 $7,032 Variable Cost $73 $0 $176 $3,418 $342 $4,009 Total Concerned that last month may not have been a typical month and her breakdown of fixed and variable costs not completely correct, the accountant decided to use the high-low method to create another cost function. She gathered individual cost and service hour data from the last six months; the following were for the two months when costs and service hours were the lowest and the highest: Cost Item Utilities Office rent Administration Supplies Training low month $222 $2,997 $3,128 $1,376 $600 $8,323 227 28 high month $321 $2,997 $3,522 $3,516 $752 $11,108 557 70 Total Service hours Customers Next month, Charlene expects the company to spend 440 hours solving clients computer system problems.REQUIRED [ROUND ALL PER-UNIT COSTS TO TNO DECIMAL PLACES AND TOTAL COSTS TO THE NEAREST DOLLAR. Part A (6 tries; 5 points) What is the accountants estimate of total overhead next month if she uses the cost function based on her analysis of last months cost data? #1 IS INCORRECT Submit Answer Incorrect. Tries 4/6 Previous Tries Part B (6 tries; 5 points) What is the accountants estimate of total overhead next month if she uses the cost function based on the high-low method? Submit Answer Incorrect. Tries 1/6 Previous Tries
Answer #1

Answer

Part A

Total Variable Cost = 73+ 0 +176 + 3418+ 342

   = 4009

Variable Cost per service hour = 4009/309

= 12.97

Total Fixed Cost = 271+ 3296 +3129 + 0 + 339

= 7032

Estimate of total 'overhead' next month = Total Fixed Cost + Variable Cost per service hour* Estimate service hour next month

Estimate of total 'overhead' next month = 7032 + 12.97*440

Estimate of total 'overhead' next month = $ 12,738.8

Part B

Total Cost at high level = 11108

Total Cost at Low level = 8323

Service hour at high level = 557

Service hour at Low level = 227

Variable Cost per service hour = (11108-8323)/(557-227)

Variable Cost per service hour = 8.43

Total Fixed Cost = 11108 - 557*8.43

Total Fixed Cost = $ 6412

Estimate of total 'overhead' next month = Total Fixed Cost + Variable Cost per service hour* Estimate service hour next month

Estimate of total 'overhead' next month = 6410 + 8.43*440

Estimate of total 'overhead' next month  = $ 10120